OpenStudy (bestinthegame):

Divide using synthetic division, 2x^3-3x^2-5x+4/x-2. Write a summary statement in fraction form.

OpenStudy (bestinthegame):

Plz Help for medal

OpenStudy (evoker):

Not a big fan of synthetic prefer long division but anyways,

OpenStudy (evoker):

First setup by finding the zero of divisor and put it to the left, while writing the coefficient of your polynomial inside the dividing region

OpenStudy (evoker):



OpenStudy (bestinthegame):

sorry not familiar with synthetic could you walk me through

OpenStudy (evoker):

Sorry have to go for a bit, here is a related article http://www.purplemath.com/modules/synthdiv.htm

OpenStudy (bestinthegame):

ok thanks

OpenStudy (robtobey2):

\[\frac{2 x^3-3 x^2-5 x+4}{x-2}=2 x^2+x-3-\frac{2}{x-2} \]